Prove the following statement by induction. For all nonnegative integers nn, 3 divides n^3 +2n +3n. State the mathematical induction and show your work clearly. 


1
Expert's answer
2021-07-06T14:34:19-0400

Let "P(n)" be the proposition that for all nonnegative integers "n, 3" divides "n^3 +2n +3."

BASIS STEP: "P(0)" is true, because "P(0)=(0)^3+2(0)+3," and "3" divides "3."

INDUCTIVE STEP: For the inductive hypothesis we assume that "P(k)" holds for an arbitrary nonnegative integer "k." That is, we assume that "3" divides "k^3 +2k +3, k=0,1,2,..."

Under this assumption, it must be shown that "P(k+1)" is true, namely, that "3" divides

"(k+1)^3 +2(k+1) +3, k=0,1,2,..."

Note that


"(k+1)^3 +2(k+1) +3"

"=k^3+3k^2+3k+1+2k+2 +3"

"=(k^3+2k+3)+3(k^2+k+1)"


We know that "3" divides "3(k^2+k+1), k=0,1,2,..."

Then "3" divides "(k^3+2k+3)+3(k^2+k+1), k=0,1,2,..." under the assumption that "P(k)" is true.

This shows that "P(k+1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all nonnegative integers "n." That is, we have proven that for all nonnegative integers "n, 3" divides "n^3 +2n +3."